using the equipartition of energy concept estimate the ideal gas Cp(not Cv) in units KB for the low temperature regime where no vibrational modes are activated(if you calculate Cp=42KB, enter 42 as your answer )
1
Expert's answer
2020-07-06T15:39:58-0400

As per equipartition theorem,

The energy of a thermodynamic system in equilibrium is partitioned equally among its degrees of freedom and molar heat capcity is given as


"C_V=0.5dk_B, C_p=C_V+k_B\\C_p=0.5(d+1)k_B"

where d is the degree of freedom, for ideal gas d=3

Now,


"C_p=0.5(3+1)k_B=2k_B"

Answer: 2